The following data relate to the operations of Shilow Company, a wholesale distributor of consumer goods:

Current assets as of March 31:
Cash $

9,300

Accounts receivable $

27,200

Inventory $

50,400

Building and equipment, net $

102,000

Accounts payable $

30,300

Common stock $

150,000

Retained earnings $

8,600

  1. The gross margin is 25% of sales.

  2. Actual and budgeted sales data:

March (actual) $ 68,000
April $ 84,000
May $ 89,000
June $ 114,000
July $ 65,000
  1. Sales are 60% for cash and 40% on credit. Credit sales are collected in the month following sale. The accounts receivable at March 31 are a result of March credit sales.

  2. Each month’s ending inventory should equal 80% of the following month’s budgeted cost of goods sold.

  3. One-half of a month’s inventory purchases is paid for in the month of purchase; the other half is paid for in the following month. The accounts payable at March 31 are the result of March purchases of inventory.

  4. Monthly expenses are as follows: commissions, 12% of sales; rent, $4,100 per month; other expenses (excluding depreciation), 6% of sales. Assume that these expenses are paid monthly. Depreciation is $765 per month (includes depreciation on new assets).

  5. Equipment costing $3,300 will be purchased for cash in April.

  6. Management would like to maintain a minimum cash balance of at least $4,000 at the end of each month. The company has an agreement with a local bank that allows the company to borrow in increments of $1,000 at the beginning of each month, up to a total loan balance of $20,000. The interest rate on these loans is 1% per month and for simplicity we will assume that interest is not compounded. The company would, as far as it is able, repay the loan plus accumulated interest at the end of the quarter.

Required:

Using the preceding data:

1. Complete the schedule of expected cash collections.

Complete the schedule of expected cash collections.

Schedule of Expected Cash Collections
April May June Quarter
Cash sales $50,400
Credit sales 27,200
Total collections $77,600

2. Complete the merchandise purchases budget and the schedule of expected cash disbursements for merchandise purchases.

Merchandise Purchases Budget
April May June Quarter
Budgeted cost of goods sold $63,000 $66,750
Add desired ending merchandise inventory 53,400
Total needs 116,400
Less beginning merchandise inventory 50,400
Required purchases $66,000
Budgeted cost of goods sold for April = $84,000 sales × 75% = $63,000.
Add desired ending inventory for April = $66,750 × 80% = $53,400.
Schedule of Expected Cash Disbursements—Merchandise Purchases
April May June Quarter
March purchases $30,300 $30,300
April purchases 33,000 33,000 66,000
May purchases
June purchases
Total disbursements

3. Complete the cash budget.

4. Prepare an absorption costing income statement for the quarter ended June 30.

5. Prepare a balance sheet as of June 30.
Answer #1
1) Shilow company
Schedule of Expected cash collections
April May June Quarter
Cash sales 50400 53400 68400 172200
credit sales 27,200 33600 35600 96,400
total collections 77600 87000 104000 268600
Accounts receivable = 114000*40%= 45600
2) Merchandise purchase budget
April May June Quarter
Budgeted cost of goods sold 63000 66750 85500 215250
Add Desired ending inventory 53400 68400 39,000 39,000
total needs 116400 135150 124500 254250
less beginning inventory 50,400 53,400 68,400 50,400
Required purchases 66,000 81,750 56,100 203,850
3) Schedule of Cash disbursements-Merchandise purhcase
April May June Quarter
March purchases 30,300 30,300
April purchases 33000 33,000 66000
May purchases 40875 40,875 81750
June purchases 28050 28050
total disbursements 63,300 73875 68925 206,100
4) Cash budget
April May June Quarter
Beginning cash balance 9,300 4,080 4,085 9,300
Add Cash collectiosn 77600 87000 104000 268600
total cas h available 86,900 91,080 108,085 277,900
less cash disbursements
for inventory 63,300 73875 68925 206,100
for expenses 19220 20120 24620 63960
for equipment 3,300 0 0 3,300
total cash disbursements 85,820 93995 93545 273,360
Excess(Deficiency)of cash 1,080 -2,915 14,540 4,540
Financing:
Borrowings 3,000 7,000 0 10,000
Repayments 0 -10,000 -10,000
interest 0 -230 -230
total financing 1,000 7,000 -10230 -230
Ending cash balance 4,080 4,085 4,310 4,310
interest = 3000*1%*3= 90
7000*1%*2= 140
230
5) income statement
Sales 287000
cost of goods sold
Beginning inventor 50,400
Add purchases 203,850
goods available for sale 254,250
ending inventory 39,000 215,250
Gross margin 71,750
Selling and administrative expense
commissions 34440
rent (4100*3) 12300
Depreciation (765*3) 2295
other expenses 17220 66255
net operating 5,495
interest expense 230
net income 5,265
Balance sheet
Assets
current assets
Cash 4,310
Accounts receivable 45,600
inventory 39,000
total current assets 88,910
Building And equipment ,net 103005
total Assets 191,915
liabilities And stockholder 's Equity
Accounts payable 28,050
total current assets 28,050
Stockholder's Equity
Capital stock 150,000
Retained earnings 13,865 163,865
total liabilites & stockholders Equity 191,915